# The Critique of Critique

### v1

@@ 4 Experiments
In this section, we conduct five experiments to demonstrate the feasibility and effectiveness of MetaCritique. The Exp. I shows the feasibility of using GPT-4 to generate reference and extract AIUs via human evaluation. The Exp. II shows the capability of various LLMs to execute the precision and recall task. The Exp. III-V show the superiority of MetaCritique variants against two GPT-4 based methods introduced by Wang et al. (2023a).
@@ 4.1 Baseline
We compare multiple modern LLMs with strong instruction-following capabilities for AIU-level precision and recall task (i.e., Exp. II). The tested models include WizardLM, LLaMA-2 Chat, Vicuna, GPT 3.5, and GPT-4. Moreover, we randomly choose 100 AIUs and engage a postgraduate student (not the annotator) to perform the same task. This result can serve as an upper benchmark (named as Human). We also introduce a baseline named GPT-4 w/o ans, where we generate the reference answer and perform the precision task in one step. It aims to confirm the importance of pre-generating a reference answer.
We compare four variants of MetaCritique with two GPT-4 based methods for the Exp. III-V:
Meta-P is the precision score of MetaCritique.
Meta-R is the recall score of MetaCritique.
Meta-F1 is the F1 score of MetaCritique.
Meta-Open is the F1 score of MetaCritique powered by open-source LLMs, whereby WizardLM 70B is used for the precision task, and WizardLM 13B is used for the recall task because they beat other open-source LLMs in the Table 4.
Pairwise is to compare two hypothesis critiques via GPT-4 and pick up the better one. The prompting instruction is shown in Table 15.
Single is to generate a likert score (1-7) for a hypothesis critique via GPT-4. The prompting instruction is shown in Table 16.
| Reference Generation | ||
| Answer | Critique | |
| 6.51 | 6.56 | |
| AIUs Extraction | ||
| Hypo.h | Hypo.l | Reference |
| 6.72 | 6.57 | 6.79 |
@@ 4.2 Exp-I: Human Evaluation for Reference Generation and AIUs Extraction
We conduct a human evaluation to validate the quality of the reference answer, reference critique, and extracted AIUs that were generated by GPT-4. We ask each human annotator to rate the outcome on a 1–7 likert score. Detailed instructions for human to evaluate reference answer generation, reference critique generation, and AIUs extraction can be found in Table 12, 13, and 14, respectively. For reference quality annotation, we present as much information as possible to annotators, including the question, model-generated answer, and hypothesis critiques. Also, we allow annotators to carefully search online whenever they need help. As for AIUs quality annotation, we only provide the original critique.
@@ Results
Table 3 shows the human evaluation for the quality of GPT-4 generated answers, critiques, and AIUs. The upper half results that GPT-4 attains remarkable performance can confirm the feasibility of using GPT-4 generated outcomes as references. Moreover, the lower half results that GPT-4 delivers impressive results can justify its use to extract AIUs for MetaCritique. 111We highly suggest using GPT-4 for reference generation and AIUs extraction for MetaCritique.
@@
| Pearson | Spearman | Kendall’s Tau | ||||
| Hypo.h | Hypo.l | Hypo.h | Hypo.l | Hypo.h | Hypo.l | |
| Single | 0.508 | 0.390 | 0.503 | 0.379 | 0.396 | 0.290 |
| Meta-Open | 0.730 | 0.709 | 0.745 | 0.690 | 0.566 | 0.506 |
| Meta-P | 0.355 | 0.667 | 0.283 | 0.681 | 0.227 | 0.504 |
| Meta-R | 0.844 | 0.831 | 0.837 | 0.830 | 0.681 | 0.649 |
| Meta-F1 | 0.841 | 0.886 | 0.836 | 0.899 | 0.675 | 0.724 |
@@
We present the AIU-level accuracy results in Table 4. We find that GPT-4 outperforms all models by a large margin and achieves performance (nearly 90%) comparable to that of human, so it is reasonable to use MetaCritique powered by GPT-4 for automated evaluation of human-written critiques and LLM-generated critiques. 222It is highly suggested to implement MetaCritique via GPT-4 due to its impressive performance in aiu-level accuracy. Moreover, WizardLM stands out as the top-performing open-source model, closely rivaling GPT-3.5. Lastly, the degradation (around 3% and 6%) without the reference answer indicates that it is necessary to pre-generate a reference answer.
@@ 4.4 Exp-III: Correlation Coefficient
In this experiment, we use different methods to generate a score for hypothesis critique. We calculate correlation coefficients to measure the correlation between different scoring baselines and human judgments. Specifically, we use Pearson correlation (Lee Rodgers and Nicewander, 1988; Mukaka, 2012), Spearman correlation (Zar, 2005), and Kendall’s Tau (Kendall, 1938) as metrics. We calculate the above correlation coefficients between the outcome score and the gold F1 score. To perform a rigorous analysis, we adopt the bootstrapping method (Koehn, 2004) for significance tests.
@@ Results
In Table 5, we show the correlation between various methods with human judgments. Our MetaCritique Meta-F1 beats the Single baseline by a large margin, confirming its increased reliability. Also, we observe that Meta-P has a reduced correlation in human-written critiques, likely because humans make fewer factuality mistakes, resulting in clustered precision scores. Lastly, Meta-Open distinctly exceeds the performance of the Single baseline, demonstrating that even less advanced LLMs can surpass GPT-4 Single baseline. This indicates that our MetaCritique framework is more effective than simple GPT-4 scoring. Lastly, precision and recall scores complement each other in assessing LLM-generated critiques. Meta-P or Meta-R is slightly inferior to Meta-F1.
4.5 Exp-IV: Agreement Rate for Pairwise Comparison
@@
In this experiment, we use various methods to determine the superior critique out of two that are generated by LLMs. We calculate the agreement rate of different methods consistent with the gold label, which stems from the gold F1 score.
@@ Results
Figure 3 illustrates the agreement rate of various methods. Our MetaCritique Meta-F1 achieves the best performance. Especially, our MetaCritique exceeds the Single baseline by a considerable margin (44%). It implies that the scores from our MetaCritique are more comparable than Single baseline. Even Meta-Open significantly outperforms GPT-4 powered baselines Single and Pairwise, confirming the effectiveness of our framework. Finally, precision and recall scores serve as complementary metrics for evaluating critiques, because Meta-P and Meta-R are somewhat less effective than Meta-F1.
@@ 4.6 Exp V: Better Critique, Better Refinement
Critique is often applied to enhance the quality of model outputs via refinement. It is intuitive that superior critiques result in better refinements. To confirm this hypothesis, we conduct this experiment. Specifically, we instruct GPT-4 to refine the model outputs via the critique. Detailed instructions with demonstrations are presented in Table 17. Subsequently, we compare the refined outcomes to choose the better one. We conduct GPT-4 evaluation and human evaluation for this comparison. The prompt for GPT-4 evaluation is shown in Table 18, while the equivalent instruction for human evaluators is outlined in Table 19.
@@ Results
We report the win rate of refined outcomes from superior critique over inferior critique, where the critique is evaluated by different methods. The outcomes of GPT-4 evaluation are depicted in Figure 4, while the results of the human evaluation appear in Figure 5. The superior critique chosen by our MetaCritique enhances refinement significantly. Moreover, precision and recall scores are mutually supportive. Relying solely on one metric can lead to diminished performance.
5 MetaCritique Leaderboard
We used MetaCritique to rank the critique models, such as GPT 3.5, SelFee (13B), UltraCM (13B), AUTO-J (13B), and human annotators stemming from Shepherd dataset. Evaluating each model for 100 critiques incurs a cost of approximately $50 when utilizing the GPT-4 API.
The results are shown in Table 6. AUTO-J is the best open-source critique model, delivering more factual and comprehensive feedback than its open-source counterparts. Also, it shows that AUTO-J surpasses other critique models in both MetaCritique recall and F1 score. In contrast, human and GPT-3.5 achieve precision scores exceeding 80%, outshining the performance of open-source critique models. This finding highlights that the research of open-source critique models should pay more attention to factuality issues.
| MetaCritique | |||
| Models | Meta-P | Meta-R | Meta-F1 |
@@ 6.1 Critique Evaluation
In light of the rapid advancements in LLMs, the significance of generating critique is increasingly acknowledged by researchers (Saunders et al., 2022; Wang et al., 2023b; Gou et al., 2023; Madaan et al., 2023; Kim et al., 2023; Ye et al., 2023; Wang et al., 2023a; Welleck et al., 2023; Li et al., 2023; Ke et al., 2023; Cui et al., 2023). However, the field is hindered by the scarcity of adequate research on critique evaluation, a gap that is currently constraining the evolution of this domain. Critique evaluation aims to evaluate the quality of the critique. It mainly relied on the human annotators (Saunders et al., 2022; Wang et al., 2023a), which entails considerable costs and carries a substantial risk of subjectivity. Wang et al. (2023a) used GPT-4 to replace human annotators, but it still lacks transparency since the numeric score is produced directly via GPT-4 without any fine-grained calculation explanation.
@@ 6.2 Meta Evaluation
Meta evaluation is designed to assess automated metrics by determining the degree of correlation between automated scores and human evaluations (Zhang* et al., 2020; Yuan et al., 2021; Sai et al., 2022; Fu et al., 2023). This is achieved by utilizing correlation coefficients. The Spearman correlation (Zar, 2005) evaluates the monotonic relationship between two variables, focusing on their ranked values rather than raw data. Conversely, the Pearson correlation (Lee Rodgers and Nicewander, 1988; Mukaka, 2012) is used to gauge the linear relationship between two variables, employing the actual data values. Furthermore, Kendall’s Tau (Kendall, 1938) is utilized to ascertain the ordinal association between two quantified variables. Besides, the significance test (Williams., 1959; Koehn, 2004) served as a crucial supplementary technique to measure the improved correlations. Lastly, accuracy sometimes serves as meta evaluation to measure the percentage of correct ranking between two outcomes (Yuan et al., 2021).
Meta evaluation is also devised to evaluate the robustness of metrics (Zhang* et al., 2020; Sai et al., 2022). Traditional metrics, which primarily rely on lexical attributes such as word overlap, can be easily attacked by text manipulation. These evaluation methods focused on adversarial scenarios or corner cases in which they fail (Callison-Burch et al., 2006; Hodosh and Hockenmaier, 2016; Kilickaya et al., 2017; Li et al., 2017; Sai et al., 2019). Recently, Zhang* et al. (2020) evaluated the robustness by using the task of adversarial paraphrase classification. The advent of generative AI has greatly mitigated this problem. Thus, we will not pursue this line of research.
@@ 6.3 Factuality Detection
Factuality detection aims to classify whether a textual statement, termed claim, is factual (Wang, 2017; Thorne et al., 2018; Augenstein et al., 2019; Zhong et al., 2020; Wadden et al., 2020; Guo et al., 2022). The FEVER dataset, introduced by Thorne et al. (2018), has led to the development of fact-checking models (Zhong et al., 2020; Krishna et al., 2022). In this context, only the claims from Wikipedia are provided to verify the claim without related evidence. Besides, factuality detection can also be formalized as factual consistency in summarization tasks, such as FactCC (Kryscinski et al., 2020) and QAGS-based models (Wang et al., 2020), which determines whether the produced summaries or summary sentences align factually with the provided evidence. In this task setting, the claim and its corresponding evidence are given. Similarly, WICEbased methods (Kamoi et al., 2023) aim to classify whether a given claim from Wikipedia can be entailed by the provided evidence. Lastly, Gao et al. (2023) and Chern et al. (2023) detected factuality without the claim and evidence. RARR (Gao et al., 2023) prompted LLMs to generate claims, retrieve evidence, and verify the claim. FacTool (Chern et al., 2023) utilized LLMs with different tools to detect factuality in various scenarios.
@@ 7 Conclusion
We are the pioneers in prioritizing critique evaluation and introducing the critique of critique, termed MetaCritique. We introduce two principles of precision and recall for the critique evaluation. Our MetaCritique is quantitative and interpretable, wherein critiques are decomposed into AIUs, and two pragmatic tasks are established to calculate objective numeric scores along with plausible natural language rationale. We collected a meta-evaluation dataset involving question answering, reasoning, entailment, and summarization to confirm the feasibility and effectiveness of our MetaCritique. Experiments also demonstrate that superior critiques chosen by our MetaCritique can result in a better refined outcome, suggesting that our MetaCritique holds the potential to substantially enhance generative artificial intelligence.
8 Limitation
The creative tasks are not suitable for recall principle, especially when there are multiple high-quality answers. It is intuitive that the suggestions for these tasks are various, so the critiques are also diverse. As a result, it is very hard to collect all effective critiques for recall principle.
Additionally, reference answers or reference critiques are often unavailable. Within our setup, GPT-4 serves as a reference, yet it is important to acknowledge that GPT-4 may commit errors.
@@ 8 Limitation
Appendix A Prompts and Instructions
We detail the prompts for GPT-4 generation, including generating hypothesis/reference critique, extracting AIUs, performing precision/recasll task, refining model-generated answer, implementing GPT-4 baselines (Single and Pairwise), and evaluating refined answer. Additionally, we provide guidelines for human evaluation, covering generated outcomes such as AIUs, reference answer/critique, refined answer. For creating reference answers, the system prompt employed in our paper is “You are a brilliant AI assistant.” We used “claims” instead of “AIUs” in our prompts to reduce possible ambiguity and confusion.

### v2

@@ 4 Experiments
In this section, we introduce various baselines and experiments to show the feasibility (Exp. I-II) and effectiveness (Exp. III-V) of MetaCritique.
@@ 4.1 Baseline
We compare multiple modern LLMs for AIU-level precision and recall tasks (Exp. II). The tested models include Zephyr (Tunstall et al., 2023), WizardLM, LLaMA-2 Chat, Vicuna, GPT 3.5, and GPT-4. Moreover, we randomly choose 100 AIUs and engage a postgraduate student (not the annotator) to perform the same task. This result can approximate a ceil performance (named as Human). We also introduce GPT-4 w/o ans, where we generate the reference answer and perform the precision task in one step. It aims to confirm the importance of pre-generating a reference answer.
We compare four variants of MetaCritique with two GPT-4 based methods introduced by Wang et al. (2023a) for Exp. III-V: MetaCritique-P, MetaCritique-R and MetaCritique-F1 is respectively the precision score, recall score and F1 score of MetaCritique powered by GPT-4. MetaCritique-F1 is the F1 score of MetaCritique powered by open-source LLMs, whereby WizardLM 70B is used for the precision task, and WizardLM 13B is used for the recall task because they beat other open-source LLMs in the Table 3. Pairwise is to compare two hypothesis critiques via GPT-4 and pick up the better one. The prompting instruction is shown in Table 16. Single is to generate a likert score (1-7) for a hypothesis critique via GPT-4. The prompting instruction is shown in Table 17.
| Reference Generation | AIUs Extraction | |||
| Answer | Critique | Hypo.h | Hypo.l | Reference |
| 6.51 | 6.56 | 6.72 | 6.57 | 6.79 |
@@ 4.2 Exp-I: Human Evaluation for Reference Generation and AIUs Extraction
We conduct a human evaluation to validate the quality of the reference answer, reference critique, and extracted AIUs that are generated by GPT-4. We engaged two postgraduate students as annotators. We ask each human annotator to rate the outcome on a 1–7 likert score. Detailed instructions for human annotators to evaluate reference answer generation, reference critique generation, and AIUs extraction can be found in Table 13, 14, and 15, respectively. To obtain a reliable evaluation outcome, we present as much information as possible to annotators. We also allow annotators to carefully search online whenever they need help.
@@ Results
Table 2 shows the rating scores for the quality of GPT-4 generated answers, critiques, and AIUs. The left half shows that GPT-4 attains remarkable performance, which can confirm the feasibility of using GPT-4 generated outcomes as references. Moreover, the right half shows that GPT-4 delivers impressive results, which can justify its effective use to extract AIUs for MetaCritique.
@@
We present the AIU-level accuracy results in Table 3. We find that GPT-4 outperforms all LLMs by a large margin and achieves an impressive performance (nearly 90%), which is comparable to that of humans. This shows that it is reasonable to use MetaCritique powered by GPT-4 for automated evaluation of human-written critiques and LLM-generated critiques. Moreover, WizardLM-70B and WizardLM-13B stand out as the top-performing open-source models in precision tasks and recall tasks, respectively. Remarkably, they rival closely GPT-3.5. Lastly, the degradation (around 3% and 6%) without the reference answer indicates that it is necessary to pre-generate a reference answer.
| Methods | Pearson | Spearman | Kendall’s Tau | |||
| Hypo.h | Hypo.l | Hypo.h | Hypo.l | Hypo.h | Hypo.l | |
| Single | 0.508 | 0.390 | 0.503 | 0.379 | 0.396 | 0.290 |
| MetaCritique-F1 | 0.730 | 0.709 | 0.745 | 0.690 | 0.566 | 0.506 |
| MetaCritique-P | 0.355 | 0.667 | 0.283 | 0.681 | 0.227 | 0.504 |
| MetaCritique-R | 0.844 | 0.831 | 0.837 | 0.830 | 0.681 | 0.649 |
| MetaCritique-F1 | 0.841 | 0.886 | 0.836 | 0.899 | 0.675 | 0.724 |
@@ 4.4 Exp-III: Correlation Coefficient
Q3: Which evaluation methods can give rating scores that are close to human judgments?
@@
In this experiment, we use different methods to generate a score for the hypothesis critique. We calculate correlation coefficients to measure the correlation between different scoring baselines and human judgments. Specifically, we use Pearson correlation (Lee Rodgers and Nicewander, 1988; Mukaka, 2012), Spearman correlation (Zar, 2005), and Kendall’s Tau (Kendall, 1938) as metrics. We calculate the above correlation coefficients between the outcome score and the gold F1 score. To perform a rigorous analysis, we adopt the bootstrapping method (Koehn, 2004) for significance tests.
@@ Results
In Table 4, we show the correlation between various methods with human judgments. Our MetaCritique-F1 beats the Single baseline by a large margin, confirming its increased reliability. We also observe that MetaCritique-P has a reduced correlation in human-written critiques, likely because humans make fewer factuality mistakes, resulting in clustered precision scores. In addition, MetaCritique distinctly exceeds the performance of the Single baseline, demonstrating that even less advanced LLMs can surpass Single baseline. This indicates that our MetaCritique framework is more effective than simple GPT-4 scoring. Lastly, precision and recall scores complement each other in assessing LLM-generated critiques. MetaCritique-P or MetaCritique-R is slightly inferior to MetaCritique-F1.
4.5 Exp-IV: Pairwise Comparison
Q4: Which evaluation methods can choose the critiques that humans prefer?
@@ 4.5 Exp-IV: Pairwise Comparison
In this experiment, we utilize a range of scoring baselines to identify the better critique out of two critiques generated by LLMs. We calculate the agreement rate to evaluate the performance, defined as the consistency of the superior critique chosen by various methods with the gold standard critique determined by gold F1 score.
@@ Results
Figure 3 illustrates the agreement rate of various methods. Our MetaCritique-F1 achieves the best performance. Especially, it exceeds the Single baseline by a considerable margin (44%). It implies that the scores of MetaCritique-F1 are more comparable than Single baseline. Even MetaCritique significantly outperforms GPT-4 powered baselines Single and Pairwise, confirming the effectiveness of our framework. Finally, precision and recall scores serve as complementary metrics for evaluating critiques, because MetaCritique-P and MetaCritique-R are somewhat less effective than MetaCritique-F1.
@@ 4.6 Exp V: Better Critique, Better Refinement
Q5: Can critique evaluations improve the alignment of existing LLMs?
@@
Critique is commonly applied to improve the quality of model outputs via refinement (Madaan et al., 2023). It is intuitive that superior critiques result in better refinements. To confirm this hypothesis, we conduct this experiment. Specifically, we instruct GPT-4 to refine the model outputs via the critique. Detailed instructions with demonstrations are presented in Table 18. Subsequently, we compare the refined outcomes to choose the better one. We conduct GPT-4 evaluation and human evaluation for this comparison. The prompt for GPT-4 evaluation is shown in Table 19, while the equivalent instruction for human evaluators is outlined in Table 20.
@@ Results
We report the win rates of refined outcomes from superior critique over inferior critique, where the critique is evaluated by different methods. The outcomes of GPT-4 evaluation are depicted in Figure 4a, while the results of the human evaluation appear in Figure 4b. The superior critique chosen by MetaCritique-F1 enhances the refinement significantly. Moreover, precision and recall scores are mutually supportive. Relying solely on one metric can lead to diminished performance.
5 Exp VI: MetaCritique Leaderboard
Q6: How do various critique models perform?
We use MetaCritique (MetaCritique-F1) to rank the critique models, such as GPT 3.5, SelFee (13B), UltraCM (13B), AUTO-J (13B), and Human annotators stemming from Shepherd dataset. As shown in Table 5, AUTO-J is the best open-source critique model, delivering more factual and comprehensive feedback than its open-source counterparts. In addition, human and GPT-3.5 achieve precision scores exceeding 80%, outshining the performance of open-source critique models. This finding highlights that the research of open-source critique models should pay more attention to factuality issues.
| Models | MetaCritique | ||
| Precision | Recall | F1 Score | |
@@ 6.1 Critique Evaluation
In light of the rapid advancements in LLMs, the significance of generating critique is increasingly acknowledged by researchers (Saunders et al., 2022; Wang et al., 2023b; Gou et al., 2023; Madaan et al., 2023; Kim et al., 2023; Ye et al., 2023; Wang et al., 2023a; Welleck et al., 2023; Li et al., 2023; Ke et al., 2023; Cui et al., 2023). However, the field is hindered by the scarcity of adequate research on critique evaluation. Critique evaluation aims to evaluate the quality of the critique. It mainly relied on the human annotators (Saunders et al., 2022; Wang et al., 2023a), which entails considerable costs and carries a substantial risk of subjectivity. Wang et al. (2023a) use GPT-4 to replace human annotators, but it still lacks transparency since the numeric score is produced directly via GPT-4 without any fine-grained calculation explanation.
@@ 6.2 Meta Evaluation
Meta evaluation is designed to assess automated metrics by determining the degree of correlation between automated scores and human evaluations (Zhang* et al., 2020; Yuan et al., 2021; Sai et al., 2022; Fu et al., 2023). This is achieved by utilizing correlation coefficients. The Spearman correlation (Zar, 2005) evaluates the monotonic relationship between two variables, focusing on their ranked values rather than raw data. Conversely, the Pearson correlation (Lee Rodgers and Nicewander, 1988; Mukaka, 2012) is used to gauge the linear relationship between two variables, employing the actual data values. Furthermore, Kendall’s Tau (Kendall, 1938) is utilized to ascertain the ordinal association between two quantified variables. Lastly, the significance test (Williams., 1959; Koehn, 2004) serves as a crucial supplementary technique to measure the improved correlations.
@@ 6.3 Factuality Detection
Factuality detection aims to classify whether a textual statement, termed claim, is factual (Wang, 2017; Thorne et al., 2018; Augenstein et al., 2019; Wadden et al., 2020; Guo et al., 2022). Thorne et al. (2018) introduce the FEVER dataset to verify the given claim without related evidence, which leads to fact-checking models (Zhong et al., 2020; Krishna et al., 2022). Besides, Kamoi et al. (2023) classify whether a given claim can be entailed by the provided evidence. Similarly, in summarization tasks, FactCC (Kryscinski et al., 2020) and QAGS-based models (Wang et al., 2020) determine whether the produced summaries or summary sentences align factually with the provided document(s). Lastly, Gao et al. (2023) and Chern et al. (2023) utilize LLMs to detect factuality without giving any claim and evidence.
@@ 7 Conclusion
We are the pioneers in prioritizing critique evaluation and introducing the critique of critique, termed MetaCritique, involving two principles: precision and recall. Our MetaCritique is quantitative, reliable, and interpretable, wherein critiques are decomposed into AIUs, and two pragmatic tasks are established to calculate objective numeric scores along with plausible natural language rationale. We curate a meta-evaluation dataset containing various tasks to confirm the feasibility and effectiveness of MetaCritique. Experiments also show that superior critiques chosen by MetaCritique can result in a better refined outcome, suggesting that MetaCritique holds the potential to enhance generative artificial intelligence. Moreover, MetaCritique can aid in identifying the pros and cons of various critique models. This paper can facilitate future research in LLM critiques.
Acknowledgements
We thank Ziqiang Cao, Yiyuan Li, and Yuqing Yang for their valuable feedback and suggestions on the writing of this paper. We also thank the anonymous reviewers for their valuable feedback and helpful suggestions. This project is supported by Qingyuan Research Project and Shanghai Artificial Intelligence Laboratory, National Natural Science Foundation of China (62076212), and Research Grants Council of Hong Kong (PolyU/15203617 and PolyU/5210919).
Limitations
The creative tasks are not suitable for the recall principle, especially when there are multiple high-quality answers. It is intuitive that the suggestions for these tasks are various, so the critiques are also diverse. As a result, it is very hard to collect all effective critiques for the recall principle.
@@ Acknowledgements
Ethical Considerations
Our experimental data stems from Shepherd, which is well-established and publicly available. Dataset construction and annotation are consistent with the intellectual property and privacy rights of the original authors. The participants in our human evaluation were transparently informed of our research intent and were paid reasonable wages. All in all, this work complies with the ACL Ethics Policy111https://www.aclweb.org/portal/content/acl-code-ethics.
Appendix A Case Study
We present a case in Figure 5, which contains more details of Figure 1, in order to illustrate how to get our MetaCritique scores.
Appendix B Gold Score
According to human annotation, we calculate the gold scores of precision, recall, and F1 score at the AIU level and critique level. The results are shown in Table 6. We find that human-written critiques get higher precision scores yet lower recall scores than LLM-generated critiques. It indicates that LLMs are even more prone to creating fictional information, while human is hardly possible to make such mistakes. Also, it is worth noting that humans miss more key information than LLMs. This might be because human-written critiques contain fewer AIUs than LLM-generated critiques.
| AIU Level (Micro) | |||
| Precision | Recall | F1 Score | |
| Hypo.h | 87.61 | 48.72 | 62.62 |
| Hypo.l | 71.85 | 53.28 | 61.19 |
| Critique Level (Macro) | |||
| Precision | Recall | F1 Score | |
| Hypo.h | 85.37 | 50.97 | 58.24 |
| Hypo.l | 71.07 | 54.37 | 58.20 |
Appendix C API Usage at PyPI
We publish MetaCritique as a Python package at PyPI in order to allow users to assess the critique conveniently. Table 7 shows how to use MetaCritique in the Python code after installation.
Appendix D Prompts and Instructions
We elaborate the prompts for GPT-4 generation, including generating hypothesis/reference critique, extracting AIUs, performing precision/recasll tasks, refining model-generated answer, implementing GPT-4 baselines (Single and Pairwise), and evaluating refined answer. Additionally, we provide guidelines for human evaluation, covering generated outcomes such as AIUs, reference answer/critique, refined answer. For creating reference answers, the system prompt employed in our paper is “You are a brilliant AI assistant.” for GPT-4. In addition, we used “claims” instead of “AIUs” in our prompts to reduce possible ambiguity and confusion for LLMs.